# Reconstruction of frequency-localized functions from pointwise samples via least squares and deep learning

### v1

@@ 1 Introduction
The study of frequency-localized functions is fundamental to signal processing and sampling theory. For example, reconstructing bandlimited functions from pointwise samples is the object of the classical Nyquist-Shannon sampling theorem, which establishes that a bandlimited signal can be reconstructed if the sampling rate is twice its bandwidth [23]. Frequency-localized and bandlimited signals are ubiquitous in engineering and signal processing, with far-reaching applications from geophysical signal estimation [39], radar imaging [12], radio transmission [11], and compressed sensing [44].
@@
Motivated by these applications, we study the approximation of multi-dimensional frequency-localized functions from uniform random samples using both least squares regression and deep learning. We start by considering least squares reconstruction methods based on the Slepian basis, or the prolate spheroidal wave functions (PSWFs) [31]. The Slepian basis, introduced by Slepian et al. at Bell Laboratories [40, 41, 42], provides an orthogonal basis for frequency-localized, bandlimited functions and has attracted attention for its unique mathematical properties. Its optimal concentration in frequency and space (or time, for univariate signals) [28, 39, 46, 48] makes it particularly effective for approximating such functions. In contrast to Chebyshev- or Lagrange-based interpolation methods, the Slepian basis has a more uniform spatial resolution and supports longer stable time steps, thus offering a distinct advantage in applications such as weather prediction [9]. Research on the Slepian basis has continued to see important developments, as highlighted in the review [47].
@@
Building on recent advances in deep learning, we further consider the approximation of smooth functions via neural networks (NNs). Here, linear expansions in the Slepian basis are replaced by nonlinear artificial NNs. This approach is motivated by the demonstrated capability of NNs to learn piecewise smooth functions [19, 36], with established dimension-independent approximation rates in the Fourier [6] and Radon domains [33]. These results strongly suggest that deep learning can provide an efficient means of modeling frequency-localized functions in applications requiring high spatial resolution.
@@
Our contribution to recovering frequency-localized functions via deep learning falls within the class of approximation theorems known as practical existence theorems (PETs), as conceived in [4] and further developed in [2]; see also [1] for a review. Traditional universal approximation theorems establish the existence of NNs capable of approximating functions from a given target class with arbitrary accuracy, provided certain conditions on network architecture are met. By leveraging results from sampling and approximation theories, PETs deliver recovery guarantees for trained NNs, contingent on sufficient conditions related to the training data and optimization strategy. Recent developments in PETs include using NNs to approximate orthogonal polynomials for basis expansions in Hilbert spaces [16, 17, 30], realizing Fourier basis functions through physics-informed NNs [10], and applying convolutional auto-encoders for reduced-order modeling [21].
@@
In this work, we prove a PET for approximating frequency-localized, bandlimited functions, using an intermediate step that represents Slepian basis functions through Legendre polynomials, which can in turn be modeled by NNs [30]. We build on results from [2], leveraging previously established findings on least squares approximation sample complexities (see, e.g., [3, Chapter 5] and references therein). This means that theoretically our NNs are assumed to be partially pre-trained. A closely related field of study to these ideas is transfer learning [32], where some layers of a previously trained model are fine-tuned on new data to solve a similar task. Specifically, this method of fine-tuning (or retraining) the last NN layer has been shown to improve the accuracy of neural networks when dealing with spurious features in the data [25, 43].
@@ 1.1 Main contributions
First, Theorem 3.1 provides a sufficient condition on the number of random samples required for least squares approximation methods to accurately recover frequency-localized functions defined over the hypercube $[-1,1]^{d}$, where $d=1,2,3$, using an expansion in Slepian functions of bandwidth $\mathsf{w}\geq 1$. The linear expansion is supported on a hyperbolic cross, a popular multi-index set used in multi-variate function approximation [18]. Specifically, we demonstrate that to recover the best approximation error in a hyperbolic cross of order $n$, it is sufficient to collect $n^{\gamma(\mathsf{w})}$ random samples (up to constants and a log factor), where $\gamma(\mathsf{w})=\lceil\log_{2}(30\mathsf{w})\rceil$.
Second, Theorem 3.2 shows that assuming the same sample complexity bound as in Theorem 3.1 it is possible to accurately approximate frequency-localized functions using deep neural networks (DNNs). In particular, this finding demonstrates the existence of a class of trainable neural networks that can be fitted to the given data by solving a least squares-type optimization problem. A key auxiliary result in our proof is Proposition 5.12, which provides a DNN construction of Slepian functions.
Third, in Section 4, we run a numerical comparison between least squares and deep learning for approximating frequency-localized functions in dimensions one and two. Our experiments show that both techniques exhibit convergence in approximation error as the number of samples increases while also underscoring the influence of the function’s domain on their performance.
@@ 2 Setup
In this section we provide the appropriate background material to present our results. Precisely, in Subsection 2.1 we provide an overview of the relevant literature on bandlimited functions, in turn leading up to the definition of Slepian functions. Then, in Subsection 2.2 we describe the framework for the least squares approximation and formally define the class of NNs considered in this paper. Before proceeding, we note that throughout this paper we denote $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ as the set of non-negative integers. We will also make use of different vector norms. Specifically, for a vector $\vec{v}=(v_{j})_{j=1}^{n}\in\mathbb{R}^{n}$, we define the $\ell^{0}$-norm and (Euclidean) $\ell^{2}$-norm of $\vec{v}$ as $\|\vec{v}\|_{0}:=\texttt{\#}\{j:v_{j}\neq 0\}$ and $\|\vec{v}\|_{2}:=(\sum_{j=1}^{n}v_{j}^{2})^{1/2}$, respectively. Moreover, by treating a matrix $A\in\mathbb{R}^{m\times n}$ as a vector, we write $\|A\|_{2}$ to denote its Frobenius norm.
@@ 2.1 Bandlimited and Slepian functions
When $d=1$ the space of $\mathsf{w}$-bandlimited functions is referred to as the Paley-Wiener space [49, Subsection 2.2], denoted by
@@
The uncertainty principle asserts that no nontrivial $f\in L^{2}(\mathbb{R})$ can have both the sets $\{x:f(x)\neq 0\}$ and $\{v:\hat{f}(v)\neq 0\}$ confined to finite measures. Nevertheless, the “Bell Labs theory” presents a framework for obtaining functions that achieve optimal concentration in both the time and frequency domains, as we now describe in greater detail.
@@
where $\mathbbm{1}_{A}$ denotes the indicator function for a set $A\subset\mathbb{R}$. It is easily verified that both $P_{\mathsf{w}}$ and $Q_{T}$ are compact self-adjoint operators on $L^{2}(\mathbb{R})$, and moreover, the range of $P_{\mathsf{w}}$ is exactly the Payley-Wiener space, i.e. $P_{\mathsf{w}}L^{2}(\mathbb{R})={\rm PW}_{\mathsf{w}}$. Since no nontrivial $f\in L^{2}(\mathbb{R})$ remains invariant under both $P_{\mathsf{w}}$ and $Q_{T}$, an alternative approach is to assess the extent to which the energy of a function in ${\rm PW}_{\mathsf{w}}$ is concentrated on $[-T,T]$. Particularly, we consider the operator $P_{\mathsf{w}}Q_{T}$, that is,
@@
A quick calculation shows that $P_{\mathsf{w}}Q_{T}f=P_{T\mathsf{w}}Q_{1}f(T\cdot)$ for all $f\in L^{2}(\mathbb{R})$, and hence without loss of generality we may set $T=1$ and simply define $\mathcal{Q}_{\mathsf{w}}:=P_{\mathsf{w}}Q_{1}$. From a straightforward calculation, we can equivalently write
@@
We are particularly interested in the eigenfunctions of $\mathcal{Q}_{\mathsf{w}}$, i.e. the solutions to the homogeneous Fredholm integral equation of the second kind, $\mathcal{Q}_{\mathsf{w}}f=\mu f$, on ${\rm PW}_{\mathsf{w}}$. It follows that the corresponding eigenvalue roughly indicates “how much” of an eigenfunction is fixed by $\mathcal{Q}_{\mathsf{w}}$. The compactness of $P_{\mathsf{w}}$ and $Q_{1}$ ensures that the spectrum of $\mathcal{Q}_{\mathsf{w}}$ is countable and accumulates at 00. We denote these eigenvalues to be
@@
where $\lim_{j\to\infty}\mu_{\mathsf{w},j}=0$ [31, Subsection 3.2.1]. An orthogonal basis for the space $L^{2}_{\bf u}([-1,1])$, which is the $L^{2}$-Lebesgue space on $[-1,1]$ with norm calibrated to the uniform measure ${\bf u}$ on $[-1,1]$, can be generated using the eigenfunctions of $\mathcal{Q}_{\mathsf{w}}$. Specifically, for $j\geq 0$, let $\varphi_{\mathsf{w},j}\in{\rm PW}_{\mathsf{w}}\cap L_{{\bf u}}^{2}([-1,1])$ to be the normalized eigenfunction of $\mu_{\mathsf{w},j}$, i.e. $\mathcal{Q}_{\mathsf{w}}\varphi_{\mathsf{w},j}=\mu_{\mathsf{w},j}\varphi_{\mathsf{w},j}$ and
@@
We refer to the set $\{\varphi_{\mathsf{w},j}\}_{j\in\mathbb{N}_{0}}$ of normalized eigenfunctions, the Slepian basis functions (the prolate spheroidal wave functions), illustrated in Figure 1 for $\mathsf{w}=16$. From (2.4), it is apparent that the function in ${\rm PW}_{\mathsf{w}}$ that minimizes the energy loss most effectively when first time-limited to $[-1,1]$ and subsequently bandlimited to $[-\mathsf{w},\mathsf{w}]$, is $\varphi_{\mathsf{w},0}$. Thus, heuristically, a function is interpreted as being $(1,\mathsf{w})$-time-frequency localized if it is a finite linear combination of eigenfunctions $\varphi_{\mathsf{w},j}$ of $\mathcal{Q}_{\mathsf{w}}$ corresponding to eigenvalues $\mu_{\mathsf{w},j}$ close to $1$, or equivalently, those with low indices $j$.
@@
with $\lim_{j\to\infty}\chi_{\mathsf{w},j}=\infty$. The progression of the spectrum (2.6) will play a key role in our subsequent analysis of the $L^{\infty}$-norm of a Slepian basis function on $[-1,1]$, a prerequisite for the proofs of the main results.
@@
for a multi-index $\vec{\nu}=(\nu_{1},\dots,\nu_{d})\in\mathbb{N}^{d}_{0}$. Continuing to denote ${\bf u}$ to be the uniform measure on $[-1,1]^{d}$, it follows from the preceding discussion that $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\mathbb{N}_{0}^{d}}$ forms an orthonormal basis of $L^{2}_{{\bf u}}([-1,1]^{d})$, which we simply refer to as the Slepian basis in $d$ dimensions.
@@ 2.2 Problem setting
To further initialize the work in this paper, in this subsection we present the main elements of our problem setting. To remain consistent with the relevant literature, the notation adopted here is analogous to that of [3, Section 5]. Letting $m\in\mathbb{N}$, we seek to approximate a continuous function $f\in\mathcal{C}([-1,1]^{d})$ from noisy samples
@@
First, we approximate $f$ using a least squares fit estimator from a hypothesis set composed of linear combinations of Slepian basis functions. Let $\Lambda\subset\mathbb{N}_{0}^{d}$ be a finite index set, whose cardinality is denoted as #$\texttt{\#}\Lambda$. Then, for a fixed $\mathsf{w}>0$, we consider the orthonormal Slepian system $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$ whose linear span $\mathcal{S}_{\Lambda}$ constitutes a (#$(\texttt{\#}\Lambda)$-dimensional subspace in $L_{{\bf u}}^{2}([-1,1]^{d})$. Supposing $m\geq\texttt{\#}\Lambda$, we approximate $f$ through $f^{\natural}$, the least squares fit among $\mathcal{S}_{\Lambda}$, given by
@@
To maintain a compact presentation, we omit the dimension $d$ in the notation $\Lambda^{\rm HC}_{n-1}$, allowing it to be inferred from the context. Continuing, we highlight two key properties of a hyperbolic cross index set. From [3, Lemma 5.15], we have the following slicing property:
@@
#$\texttt{\#}\Lambda_{j}\leq\texttt{\#}\Lambda_{k}.$ (2.10)
Second, we establish the existence of a class of NNs capable of Slepian basis functions. For this purpose, we introduce the formal definition of an NN, as follows.
Definition 2.3 ([34, Definition 2.1]).

### v2

@@ 1 Introduction
The study of frequency-localized functions is fundamental to signal processing and sampling theory. For example, reconstructing bandlimited functions from pointwise samples is the object of the classical Nyquist-Shannon sampling theorem, which establishes that a bandlimited signal can be reconstructed if the sampling rate is twice its bandwidth [jerri1977shannon]. Frequency-localized and bandlimited signals are ubiquitous in engineering and signal processing, with far-reaching applications from geophysical signal estimation [simons2010slepian], radar imaging [chen2002highly], radio transmission [cabric2004implementation], and compressed sensing [tropp2009beyond].
@@
Motivated by these applications, we study the approximation of multi-dimensional frequency-localized functions from uniform random samples using both least squares regression and deep learning. Specifically, we consider least squares reconstruction methods based on the Slepian basis, or the prolate spheroidal wave functions (PSWFs) [osipov2013prolate]. Introduced by Slepian et al. at Bell Laboratories [slepian1961prolate, slepian1964prolate, slepian1965some], this orthogonal basis consists of functions optimally concentrated in both space (or time) and frequency [moore2004prolate, simons2010slepian, walter2003sampling, xiao2001prolate], making it particularly effective for approximating frequency-localized or bandlimited functions confined to a bounded spatial domain. Continued research on the Slepian basis has underscored its distinctive mathematical properties, as reviewed in [wang2017review]. We include a brief discussion on the numerical benefits of using Slepian functions in applications at the end of Section 2.1.
@@
Building on recent advances in deep learning, we further consider the approximation of smooth functions via neural networks (NNs). Here, linear expansions in the Slepian basis are replaced by nonlinear artificial NNs. This approach is motivated by the demonstrated capability of NNs to learn piecewise smooth functions [eckle2019comparison, petersen2018optimal], with established dimension-independent approximation rates in the Fourier [barron1993universal] and Radon domains [parhi2022near]. These results strongly suggest that deep learning can provide an efficient means of modeling frequency-localized functions in applications requiring high spatial resolution.
@@
Our contribution to recovering frequency-localized functions via deep learning falls within the class of approximation theorems known as practical existence theorems (PETs), as conceived in [adcock2021gap] and further developed in [adcock2022deep]; see also [adcock2024learning] for a review. Traditional universal approximation theorems establish the existence of NNs capable of approximating functions from a given target class with arbitrary accuracy, provided certain conditions on network architecture are met. By leveraging results from sampling and approximation theories, PETs deliver recovery guarantees for trained NNs, contingent on sufficient conditions related to the training data and optimization strategy. Recent developments in PETs include using NNs to approximate orthogonal polynomials for basis expansions in Hilbert spaces [daws2019analysis, de2021approximation, opschoor2022exponential], realizing spectral Fourier collocation approximations to PDE solutions through physics-informed NNs [brugiapaglia2024physics], and applying convolutional auto-encoders for reduced-order modeling [franco2025practical].
@@
In this work, we prove a PET for approximating frequency-localized, bandlimited functions, using an intermediate step that represents Slepian basis functions through Legendre polynomials, which can in turn be modeled by NNs [opschoor2022exponential]. We build on results from [adcock2022deep], leveraging previously established findings on least squares approximation sample complexities (see, e.g., [adcock2022sparse, Chapter 5] and references therein). This means that theoretically our NNs are assumed to be partially pre-trained. A closely related field of study to these ideas is transfer learning [pan2009survey], where some layers of a previously trained model are fine-tuned on new data to solve a similar task. Specifically, this method of fine-tuning (or retraining) the last NN layer has been shown to improve the accuracy of neural networks when dealing with spurious features in the data [kirichenko2022last, strombergenhancing].
@@ 1.1 Main contributions
First, Theorem 3.1 provides a sufficient condition on the number of random samples required for least squares approximation methods to accurately recover frequency-localized functions defined over the hypercube $1,1]^{d}$, where $d=1,2,3$, using an expansion in Slepian functions of bandwidth $\mathsf{w}\geq 1$. The linear expansion is supported on a hyperbolic cross, a popular multi-index set used in multi-variate function approximation [dung2018hyperbolic]. Specifically, we demonstrate that to recover the best approximation error in a hyperbolic cross of order $n$, it is sufficient to collect $n^{\gamma(\mathsf{w})}$ random samples, up to a constant and a logarithmic factor, where $\gamma(\mathsf{w})=\lceil\log_{2}(24\mathsf{w})\rceil$.
Second, Theorem 3.3 shows that assuming the same sample complexity bound as in Theorem 3.1 it is possible to accurately approximate frequency-localized functions using deep neural networks (NNs). In particular, this finding demonstrates the existence of a class of trainable neural networks that can be fitted to the given data by solving a least squares-type optimization problem. A key auxiliary result in our proof is Proposition 5.10, which provides a NN construction of Slepian functions.
Third, in Section 4, we run a numerical comparison between least squares and deep learning for approximating frequency-localized functions in dimensions one and two. Our experiments show that both techniques exhibit convergence in approximation error as the number of samples increases while also underscoring the influence of the function’s domain on their performance. Moreover, we propose a Slepian-based initialization strategy inspired by the proof of Theorem 3.3 and show that it can outperform standard random initializations in NN training.
@@ 2 Setup
In this section we provide the appropriate background material to present our results. Precisely, in Section 2.1 we provide an overview of the relevant literature on bandlimited functions, in turn leading up to the definition of Slepian functions. Then, in Section 2.2 we describe the framework for the least squares approximation and formally define the class of NNs considered in this paper. Before proceeding, we note that throughout this paper we denote $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ as the set of non-negative integers. For a vector $\vec{v}=(v_{j})_{j=1}^{n}\in\mathbb{C}^{n}$, we will make use of the (Euclidean) $\ell^{2}$-norm and the so-called $\ell^{0}$-norm of $\vec{v}$, defined as $\|\vec{v}\|_{2}:=(\sum_{j=1}^{n}|v_{j}|^{2})^{1/2}$ and $\|\vec{v}\|_{0}:=\texttt{\#}\{j:v_{j}\neq 0\}$, respectively. Moreover, by treating a matrix $A\in\mathbb{R}^{m\times n}$ as a vector, we write $\|A\|_{2}$ to denote its Frobenius norm.
@@ 2.1 Bandlimited and Slepian functions
When $d=1$ the space of $\mathsf{w}$-bandlimited functions is referred to as the Paley-Wiener space [zayed2018advances, Section 2.2], denoted by
@@
The uncertainty principle asserts that no nontrivial $f\in L^{2}(\mathbb{R})$ can have both the set $\{x:f(x)\neq 0\}$ and the set $\{v:\hat{f}(v)\neq 0\}$ confined to finite measures (see, e.g., [amrein1977support, Proposition 8] or [benedicks1985fourier, Theorem 2]). Nevertheless, the “Bell Labs theory” presents a framework for obtaining functions that achieve optimal concentration in both the time and frequency domains, as we now describe in greater detail.
@@
where $\mathbbm{1}_{A}$ denotes the indicator function for a set $A\subset\mathbb{R}$. It is readily verified that $P_{\mathsf{w}}(L^{2}(\mathbb{R}))={\rm PW}_{\mathsf{w}}$. Since no nontrivial $f\in L^{2}(\mathbb{R})$ remains invariant under both $P_{\mathsf{w}}$ and $Q_{T}$, an alternative approach is to assess the extent to which the energy of a function in ${\rm PW}_{\mathsf{w}}$ is concentrated on [T,T][-T,T]. Particularly, we consider the operator $P_{\mathsf{w}}Q_{T}$, that is,
@@
The operator $P_{\mathsf{w}}Q_{T}$ is a Hilbert-Schmidt, hence compact, operator on ${\rm PW}_{\mathsf{w}}$ (see, e.g., [hogan2005time, Section 3.3.3]). A quick calculation shows that $P_{\mathsf{w}}Q_{T}f=P_{T\mathsf{w}}Q_{1}f(T\cdot)$ for all $f\in L^{2}(\mathbb{R})$, and hence without loss of generality we may set $T=1$ and simply define $\mathcal{Q}_{\mathsf{w}}:=P_{\mathsf{w}}Q_{1}$. From a straightforward calculation, we can equivalently write
@@
We are particularly interested in the eigenfunctions of $\mathcal{Q}_{\mathsf{w}}$, i.e., the solutions to the homogeneous Fredholm integral equation of the second kind, $\mathcal{Q}_{\mathsf{w}}f=\mu f$, on ${\rm PW}_{\mathsf{w}}$. It follows that the corresponding eigenvalue roughly indicates “how much” of an eigenfunction is fixed by $\mathcal{Q}_{\mathsf{w}}$. The spectrum of $\mathcal{Q}_{\mathsf{w}}$ is countable and accumulates at 0; we denote these eigenvalues as
@@
where $\lim_{j\to\infty}\mu_{\mathsf{w},j}=0$ [osipov2013prolate, Section 3.2.1]. An orthogonal basis for the space $L^{2}_{\bf u}([-1,1])$, which is the $L^{2}$-Lebesgue space on [1,1][-1,1] with norm calibrated to the uniform measure $\bf u$} on [1,1][-1,1], can be generated using the eigenfunctions of $\mathcal{Q}_{\mathsf{w}}$. Specifically, for $j\geq 0$, let $\varphi_{\mathsf{w},j}\in{\rm PW}_{\mathsf{w}}\cap L_{{\bf u}}^{2}([-1,1])$ to be the normalized eigenfunction associated with $\mu_{\mathsf{w},j}$, i.e., $\mathcal{Q}_{\mathsf{w}}\varphi_{\mathsf{w},j}=\mu_{\mathsf{w},j}\varphi_{\mathsf{w},j}$ and
@@
We refer to the set $\{\varphi_{\mathsf{w},j}\}_{j\in\mathbb{N}_{0}}$ of normalized eigenfunctions, the Slepian basis functions (the prolate spheroidal wave functions), illustrated in Figure 1 for $\mathsf{w}=16$. From (2.4), it is apparent that the function in ${\rm PW}_{\mathsf{w}}$ that minimizes the energy loss most effectively when first time-limited to [1,1][-1,1] and subsequently bandlimited to $\mathsf{w},\mathsf{w}]$, is $\varphi_{\mathsf{w},0}$. Thus, heuristically, a function is interpreted as being $(1,\mathsf{w})$-time-frequency localized if it is a finite linear combination of eigenfunctions $\varphi_{\mathsf{w},j}$ of $\mathcal{Q}_{\mathsf{w}}$ corresponding to eigenvalues $\mu_{\mathsf{w},j}$ close to $1$, or equivalently, those with low indices $j$.
@@
with $\lim_{j\to\infty}\chi_{\mathsf{w},j}=\infty$. The progression of the spectrum (2.6) will play a key role in our subsequent analysis of the $L^{\infty}$-norm of a Slepian basis function on [1,1][-1,1], a prerequisite for the proofs of the main results.
@@
for a multi-index $\vec{\nu}=(\nu_{1},\dots,\nu_{d})\in\mathbb{N}^{d}_{0}$. Continuing to denote $\bf u$} to be the uniform measure on $1,1]^{d}$, it follows from the preceding discussion that $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\mathbb{N}_{0}^{d}}$ forms an orthonormal basis of $L^{2}_{{\bf u}}([-1,1]^{d})$, which we simply refer to as the Slepian basis in $d$ dimensions.
Remark 2.2 (Practical advantages of Slepian representations).
Slepian functions are specifically designed for applications involving spatiospectral concentration, such as spectral estimation or problems constrained by geometry. This reflects a fundamental principle in approximation theory: an approximation is most accurate and efficient when the basis mirrors the intrinsic structure of the target space. The practical utility of Slepian functions and their variants (Slepian-like bases) is well documented. For instance, [simons2006spatiospectral] employed them to obtain leakage-controlled spectral estimates of geophysical data localized to Earth’s continents and oceans, in contrast to spherical harmonics, whose energy is distributed globally.
Associated with a spatiospectral concentration problem is its Shannon number, which quantifies the effective dimensionality of the space of functions simultaneously localized in space (or time) and frequency. This number is known to be roughly proportional to the product between the spatial (or temporal) and frequency domain areas. Remarkably, a defining feature of the corresponding Slepian basis is the rapid decay of their eigenvalues (see e.g. (2.4)) beyond a certain index, which in one dimension has been rigorously shown to closely align with the Shannon number (see [karnik2021improved, Corollary 3] or (3.10)). As a result, only the first few Slepian functions contribute significantly, while the remainder have negligible impact, enabling sparse and efficient representations of similarly localized functions. This behavior naturally yields a dimensionality reduction determined entirely by the domain, to which the Slepian basis is inherently adapted. In contrast, classical bases such as Chebyshev or Legendre polynomials provide no intrinsic mechanism for such compression.
Moreover, Slepian functions offer notable advantages in spectral and spectral-element methods for nearly bandlimited, highly oscillatory solutions, primarily due to their nearly uniform oscillatory behavior, which induces quasi-uniform spatial grids. This uniformity improves spatial resolution and allows for longer stable time steps, giving the Slepian basis a distinct advantage in applications such as weather prediction. By contrast, Chebyshev and Legendre polynomials have non-uniform oscillations that significantly cluster near the boundaries, limiting their efficiency in high-resolution computations [boyd2004prolate]. Similarly, for eigenvalue problems [wang2017review, Section 5.3], Slepian-based collocation or Galerkin methods deliver superior resolution of oscillatory waves compared with traditional polynomial bases, requiring fewer points per wavelength to achieve a given accuracy.
@@ 2.2 Problem setting
To further initialize the work in this paper, in this subsection we present the main elements of our problem setting. To remain consistent with the relevant literature, the notation adopted here is analogous to that of [adcock2022sparse, Section 5]. Letting $m\in\mathbb{N}$, we seek to approximate a continuous function $f\in\mathcal{C}([-1,1]^{d})$ from noisy samples
@@
First, we approximate $f$ using a least squares fit estimator from a hypothesis set composed of linear combinations of Slepian basis functions. Let $\Lambda\subset\mathbb{N}_{0}^{d}$ be a finite index set, whose cardinality is denoted as $\texttt{\#}\Lambda$. Then, for a fixed $\mathsf{w}>0$, we consider the orthonormal Slepian system $\{\varphi_{\mathsf{w},\vec{\nu}}\}_{\vec{\nu}\in\Lambda}$ whose linear span $\mathcal{S}_{\Lambda}$ constitutes a $(\texttt{\#}\Lambda)$-dimensional subspace in $L_{{\bf u}}^{2}([-1,1]^{d})$. Supposing $m\geq\texttt{\#}\Lambda$, we approximate $f$ through $f^{\natural}$, the least squares fit among $\mathcal{S}_{\Lambda}$, given by
@@
To maintain a compact presentation, we omit the dimension $d$ in the notation $\Lambda^{\rm HC}_{n-1}$, allowing it to be inferred from the context. [^1] Among its desirable properties, the hyperbolic cross has a well-understood cardinality: $\texttt{\#}\Lambda^{\rm HC}_{n-1}\sim n(\log n)^{d-1}/(d-1)!$ for $n\to\infty$ and any fixed $d\in\mathbb{N}$ (see [dobrovol1998number]). A nonasymptotic bound can be found in Lemma 6.1. Continuing, we highlight two key properties of a hyperbolic cross index set. From [adcock2022sparse, Lemma 5.15], we have the following slicing property: [^2] Note that in this union, the level $k$ stops at n1n-1. Moreover—although we do not use this explicitly—the sets $\Lambda_{k}$ are nested, meaning $\Lambda_{j}\subset\Lambda_{k}$ if $j\geq k$. This property, in turn, implies the monotonicity property (2.10).
@@
$\texttt{\#}\Lambda_{j}\leq\texttt{\#}\Lambda_{k}.$ (2.10)
Second, we establish the existence of a class of NNs capable of approximating the Slepian basis functions. To this end, we introduce the formal definition of an NN, as follows.
Definition 2.4 ([PetV2018OptApproxReLU, Definition 2.1]).